Hello,
my daughter and I are working on a few math problems she brought home from school and It's been some time for me to do it without review...One one the problems is:
4/9v=1/4
we need to solve for v
is that clear enough? I hope so and thank you so much for your help and the help of this website...
Thank You,
Seann

Answer by stanbon(75887)

(4/9)v=1/4
we need to solve for v
------------------------
Multiply both sides by (9/4):
(9/4)(4/9)v = (9/4)(1/4)
Cancel numbers that are common to a numerator and a denominator.
v = 9/16
=====================
Cheers,
Stan H.